# Structure-Preserving Model Order Reduction for Nonlinear DAE Models of Power Networks

### v1

@@ Abstract
This paper deals with the joint reduction of dynamic states (internal states of generator, solar, and loads, etc) and algebraic variables (states of the network e.g., voltage and phase angles) of a nonlinear differential-algebraic equation (NDAE) model of power networks. Traditionally, in the current literature of power system model order reduction (MOR), the algebraic constraints are usually neglected and the power network is commonly modeled via a set of ordinary differential equations (ODEs) instead of NDAEs. Thus, reduction is usually carried out for the dynamic states only and the algebraic variables are kept intact. This leaves a significant part of the system’s size and complexity unreduced. This paper addresses this aforementioned limitation, by jointly reducing both dynamic and algebraic variables. As compared to the literature the proposed MOR techniques herein are endowed with the following features: (i) no system linearization is required, (ii) requires no transformation to an equivalent or approximate ODE representation, (iii) guarantee that the reduced order model to be NDAE and thus preserves the differential-algebraic structure of original power system model, and (iv) can seamlessly reduce both dynamic and algebraic variables while maintaining high accuracy. Case studies performed on a 2000-bus power system reveal that the proposed MOR techniques are able to reduce system order while maintaining accuracy.
@@ I Introduction and Paper Contributions
[lines=2]Power systems form large-scale, complex networks that usually require large state-space expressions for accurate modeling. The complexity and size of power systems are even further increasing with the integration of renewables and other (power-electronics)-based distributed resources. Thus, the research area of model order reduction (MOR) in power systems is becoming highly crucial. Specifically, for the design of advanced feedback controllers (e.g., $\mathcal{H}_{2}$, $\mathcal{H}_{\infty}$, and LQR/LQG), the design of reduced-order model (ROM) is of extreme importance. This is because the order of these controllers matches the order of the system, thus for a very large system with thousands of state variables, the design of these controllers becomes intractable [1, 2].
The idea of model reduction is not new in power systems research, and significant relevant research has been proposed in the past two decades. Generally speaking, there are two main MOR philosophies in the current literature of power systems. The first philosophy divides the power system into study (internal) and external areas. The study area is represented in detail, while the external area is simplified or approximated. This approach leverages coherency-based methods [3, 4], which are rooted in the identification of coherent generator groups within the power system. The reduction process generally involves three steps: $(i)$ identifying coherency among generators, $(ii)$ dynamically reducing the system by aggregating the network and generators, and $(iii)$ potentially aggregating excitation controllers also in later stages. These coherency-based methods are highly regarded for their reliability in achieving dynamic equivalence in power systems. However, a notable limitation is the potential inability to reduce specific parts of the power network due to the inherent nature of coherency grouping [5].
@@
These MOR methodologies have also been widely applied to power systems to construct various ROMs. For instance, authors in [8, 1] have proposed moment matching-based MOR techniques. In [9, 10], researchers have proposed balanced realization-based MOR algorithms for linear ODE-based power system models. Later, these works have been extended in [5] to propose MOR for nonlinear ODE-based power system models where instead of using Grammians, empirical controllability and observability covariance matrices are used to balance and truncate the system. Similarly, in [11, 12] various modal truncation-based reduction algorithms have been proposed. Readers are referred to [13, 14] for a detailed survey of the existing MOR approaches in power systems.
However, in most of the current MOR power system literature, the algebraic constraints (modeling power/current balance) are usually neglected and the power system is modeled via a set of linear ODEs (or converted to linear ODEs) in order to apply MOR concepts from linear system control theory. The recent work in [5] has considered nonlinearity in their design. Yet again, the algebraic constraints are neglected and thus MOR is only carried out for dynamic variables while keeping the algebraic variables intact. This is problematic as algebraic variables often constitute a large portion of the model, representing essential electrical quantities across the network. Ignoring these in the reduction process means a significant part of the system’s complexity and size remains untouched, limiting the effectiveness of MOR. Also, considering the complete NDAE dynamics is essential because the linear/nonlinear ODE-based power system models cannot capture the effects of topological changes (like the tripping of transmission lines, etc.) [15]. Furthermore, it is unclear how to include dynamics of loads and renewables in the ODE-based models [16].
@@
  • Due to the diagonal structure of the proposed coordinate transformation matrix, the presented techniques ensure that the ROMs are NDAEs similar to the original power system model. This preserves the differential-algebraic structure of power networks. This also allows a seamless transition from reduced order to the full order dynamic and algebraic state variables.
  • Thorough time-domain simulation studies under different transient conditions have been carried out to assess the performance of the proposed techniques. The considered test system includes: (i) modified IEEE 39-bus system modeling detailed conventional power plants dynamics, solar plant acting in grid-forming mode, composite load dynamics, and algebraic constraint models, and (ii) a 2000-bus Texas network with comprehensive $11^{th}$-order conventional power plant models. Furthermore, to showcase the advantage of the proposed MOR techniques a comparison with the commonly used linear ODE-based balance truncation method has also been presented.
Paper Organization.   Sec. II presents the advanced power system model considered in this study. Sec. III delineates the problem formulation and scope. Sec. IV and V present the two MOR algorithms for NDAE power system models. Case studies are presented in Sec. VI.
Notation.   Capital bold letters are used to represent matrices, as in $\mathbf{A}$ while small capital bold letters, such as $\mathbf{b}$, denote vectors. All the sets are represented in calligraphic fonts, such as $\mathcal{M}$ or $\mathcal{R}$. The symbol $\mathbb{R}^{u\times v}$ represents a real-valued matrix of size $u\times v$, similarly, $\mathbb{R}^{k}$ denotes a real-valued column vector with $k$ elements. The notations $\bm{O}$ and $\bm{I}$ denote zero and identity matrix of appropriate dimensions, respectively. The union of two sets is denoted by $\cup$. The symbol $\mathbb{S}^{n\times n}_{++}$ represents a square positive definite matrix of size $n\times n$ while the notation $\bm{e}_{j}$ represents a column vectors of zeros with $1^{\prime}s$ only at location $j$.
II Nonlinear DAE power system models
We consider a grid model with $S$ solar power plants, $G$ steam and hydro-based conventional power plants, and $L_{p}$, $L_{k}$, $L_{z}$ number of constant power loads, motor-based, and constant impedance loads, respectively. The overall power system is modeled as a graph with $\mathcal{N}=\left\{1,...,N\right\}$ as the set of nodes/buses and $\mathcal{E}$ as the set of edges or transmission lines. The set of buses are grouped into various types: $\mathcal{S}=\left\{1,...,S\right\}$ represents buses with PV power plants, $\mathcal{G}=\left\{1,...,G\right\}$ denotes buses connected to the conventional power plants, $\mathcal{L}$ includes buses that contain constant power/impedance and motor-based loads, while $\mathcal{U}$ collects non-unit buses that are not connected to any elements.
@@ II Nonlinear DAE power system models
In model (1), the set of differential equations (1a) encompasses the dynamics of PV plants, conventional power plants, and composite loads dynamics, while the algebraic constraint (1b) models the current/power balance equations of the electrical network. The vector $\bm{x}_{d}(t)\in\mathbb{R}^{n_{d}}$ represents dynamic states and it lumps the dynamic variables of conventional power plants, PV plants, and composite load dynamics as $\bm{x}_{d}=\begin{bmatrix}\bm{x}_{G}^{\top}&\bm{x}_{R}^{\top}&\bm{x}_{L}^{\top}\end{bmatrix}^{\top}$ where $\bm{x}_{G}$ are the states of conventional power plants, $\bm{x}_{R}$ denotes the states of PV plants, and $\bm{x}_{L}$ represents the states of dynamic loads. The notation $\bm{x}_{a}(t)\in\mathbb{R}^{n_{a}}$ denotes algebraic states and it contains the states of the network (voltage and current phasors). The vector $\bm{w}(t)=\begin{bmatrix}\bm{P}_{d}^{\top}&\bm{I}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{w}}$ contains load demand $\bm{P}_{d}$ and sun irradiance $\bm{I}$ while $\bm{u}(t)=\begin{bmatrix}\bm{u}_{G}^{\top}&\bm{u}_{R}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{u}}$ defines the system control inputs with $\bm{u}_{G}$ and $\bm{u}_{R}$ denoting the control inputs of generators and solar farms, respectively. Further detailed explanations of these vectors and complete dynamical equations (set of differential equations) describing the models of PV plants, conventional power plants, composite load dynamics, and system algebraic constraint model used in this study are given in Appendix A.
@@
where $\bm{E}\in\mathbb{R}^{n\times n}$ is a singular binary matrix encoding system algebraic model with rows of zeros, function ${\bm{f}}\left({\bm{x}},{\bm{u}},{\bm{w}}\right)$ represents the corresponding nonlinearity, while the rest of the real-valued matrices ${\bm{B}}_{u}\in\mathbb{R}^{n\times n_{u}}$, ${\bm{A}}\in\mathbb{R}^{n\times n}$, ${\bm{B}}_{w}\in\mathbb{R}^{n\times n_{w}}$ maps the system control inputs $\bm{u}$, state vector $\bm{x}$, and the disturbance vector $\bm{w}$ in the power system dynamics. The output from power system model is considered to be bus voltages and current phasors which are mapped from state vector $\bm{x}$ via output matrix $\bm{C}\in\mathbb{R}^{p\times n}$. Furthermore, throughout the paper, we assume that in the NDAE model (2), the pair $(\bm{E},\bm{A})$ is regular, and the power system model is observable and controllable. These assumptions are common and power system NDAEs are known to be regular, controllable, and observable [2, 19].
@@ IV Structure-Preserving POD (SP-POD) MOR
Therefore, we reduce the nonlinearity using the discrete empirical interpolation method (DEIM)-based approach [7]. The main goal is to handle the nonlinear terms effectively in a reduced-dimensional space $\mathbb{R}^{r}$ instead of the full-dimensional space $\mathbb{R}^{n}$. This is done by measuring specific points in the state space rather than the entire set of state variables. Then, the nonlinear term is approximated by interpolating around selected points, with the number of measured points matching the reduced-dimensional space $r$. This approximation can be expressed as:
$\underbrace{\bm{W}_{L}\bm{W}_{fr}}_{r\times r}\underbrace{{\bm{f}}_{r}\left(\cdot\right)}_{r\times 1}.$ (9)
@@
To carry out such approximation for the nonlinear term we start by storing the snapshots of the ${\bm{f}}\left({{\bm{x}}},{\bm{u}},{\bm{w}}\right)$ in a matrix (by performing time-domain simulation similar to as done while designing $\bm{X}_{d}$ and $\bm{X}_{a}$ in (5)) given as:
@@
Then we take SVD of this snapshot matrix, $\bm{X}_{f}=\bm{W}_{f}\bm{\Sigma}_{f}\bm{\Lambda}_{f}$ with matrix $\bm{W}_{f}=\begin{bmatrix}\bm{w}_{f_{1}}\;,\bm{w}_{f_{2}}\;,\cdots,\bm{w}_{f_{n}}\end{bmatrix}\in\mathbb{R}^{n\times n}$ containing the left singular vectors, matrix $\bm{\Lambda}_{f}\in\mathbb{R}^{t\times n}$ lumping the right singular vectors, and $\bm{\Sigma}_{f}\in\mathbb{R}^{n\times n}$ being the diagonal matrix containing the HSVs. Next, we design $\bm{W}_{fr}=\begin{bmatrix}\bm{w}_{f_{1}}\;,\bm{w}_{f_{2}}\;,\cdots,\bm{w}_{f_{r}}\end{bmatrix}\in\mathbb{R}^{n\times r}$ as the first $r$ columns of $\bm{W}_{f}$. Finally, we design a binary measurement matrix $\bm{P}_{M}$ that selects optimal points in the reduced subspace $\bm{W}_{fr}$ so that nonlinearity can be reconstructed using the selected points efficiently. To construct such $\bm{P}_{M}$ we utilize residual-based greedy technique [7], which essentially puts a measurement (or a $1$ in matrix $\bm{P}_{M}$) where residual/error is maximum (highlighting measurement point is required). The designed greedy algorithm proceeds as follows:
@@
and the next measurement point is selected where the value of $\bm{d}$ is maximum. By selecting points with the maximum residual, the algorithm ensures that each new measurement location adds the most significant new information about the nonlinear term $\bm{f}(\cdot)$. The iterative process continues until the number of chosen measurement points equals the dimension $r$ of the reduced subspace. After completion of the iterations, the approximation to the nonlinearity can be expressed as:
@@
Hence, the final ROM can be written as follows:
@@
The overall proposed SP-POD-based MOR algorithm is summarized in Algorithm 1 given in Appendix B. In the following section, we present the structure-preserving balanced POD-based (SP-BPOD) MOR technique for the complete NDAE representation of power systems.

### v2

@@ Abstract
This paper deals with the joint reduction of the number of dynamic and algebraic states of a nonlinear differential-algebraic equation (NDAE) model of a power network. The dynamic states depict the internal states of generators, loads, renewables, whereas the algebraic ones define network states such as voltages and phase angles. In the current literature of power system model order reduction (MOR), the algebraic constraints are usually neglected and the power network is commonly modeled via a set of ordinary differential equations (ODEs) instead of NDAEs. Thus, reduction is usually carried out for the dynamic states only and the algebraic variables are kept intact. This leaves a significant part of the system’s size and complexity unreduced. This paper addresses this aforementioned limitation by jointly reducing both dynamic and algebraic variables. As compared to the literature the proposed MOR techniques are endowed with the following features: (i) no system linearization is required, (ii) require no transformation to an equivalent or approximate ODE representation, (iii) guarantee that the reduced order model to be NDAE-structured and thus preserves the differential-algebraic structure of original power system model, and (iv) can seamlessly reduce both dynamic and algebraic variables while maintaining high accuracy. Case studies performed on a 2000-bus power system reveal that the proposed MOR techniques are able to reduce system order while maintaining accuracy.
@@ I Introduction and Paper Contributions
[lines=2]Power systems form large-scale, complex networks that usually require large state-space expressions for accurate modeling. The complexity and size of power systems are even further increasing with the integration of renewables and other (power-electronics)-based distributed resources. Thus, the research area of model order reduction (MOR) in power systems is becoming highly crucial. Specifically, for the design of advanced feedback controllers (e.g., $\mathcal{H}_{2}$, $\mathcal{H}_{\infty}$, and LQR/LQG), the design of reduced-order model (ROM) is of extreme importance. This is because the order of these controllers matches the order of the system, thus for a very large system with thousands of state variables, the design of these controllers becomes intractable [1, 2, 3].
The idea of model reduction is not new in power systems research, and significant relevant research has been proposed in the past two decades. Generally speaking, there are two main MOR philosophies in the current literature of power systems. The first philosophy divides the power system into study (internal) and external areas. The study area is represented in detail, while the external area is simplified or approximated. This approach leverages coherency-based methods [4, 5, 6, 7], which are rooted in the identification of coherent generator groups within the power system. The reduction process generally involves three steps: $(i)$ identifying coherency among generators, $(ii)$ dynamically reducing the system by aggregating the network and generators, and $(iii)$ potentially aggregating excitation controllers also in later stages. These coherency-based methods are highly regarded for their reliability in achieving dynamic equivalence in power systems. However, a notable limitation is the potential inability to reduce specific parts of the power network due to the inherent nature of coherency grouping [8].
@@
These MOR methodologies have also been widely applied to power systems to construct various ROMs. For instance, authors in [11, 1] have proposed moment matching-based MOR techniques. In [12, 13, 14], researchers have proposed balanced realization-based MOR algorithms for linear ODE-based power system models. Later, these works have been extended in [8] to propose MOR for nonlinear ODE-based power system models where instead of using Grammians, empirical controllability and observability covariance matrices are used to balance and truncate the system. Similarly, in [15, 16, 17] various modal truncation-based reduction algorithms have been proposed. Moreover, a Loewner matrix method-based approach for efficient model order reduction and system identification in power systems was recently proposed in [18]. Additionally, [19] introduced a parametric MOR technique that preserves critical device parameters while effectively reducing large-scale power system models. Readers are referred to [20, 21] for a detailed survey of the existing MOR approaches in power systems.
However, in most of the current MOR power system literature, the algebraic constraints (modeling power/current balance) are usually neglected and the power system is modeled via a set of linear ODEs (or converted to linear ODEs) in order to apply MOR concepts from linear system control theory. The recent work in [8] has considered nonlinearity in their design. Yet again, the algebraic constraints are neglected and thus MOR is only carried out for dynamic variables while keeping the algebraic variables intact. This is problematic as algebraic variables often constitute a large portion of the model, representing essential electrical quantities across the network. Ignoring these in the reduction process means a significant part of the system’s complexity and size remains untouched, limiting the effectiveness of MOR. Also, considering the complete nonlinear differential-algebraic (NDAE) dynamics is essential because the linear/nonlinear ODE-based power system models cannot capture the effects of topological changes (like the tripping of transmission lines, etc.) [22].
@@
  • Due to the diagonal structure of the proposed coordinate transformation matrix, the presented techniques ensure that the ROMs are NDAEs similar to the original power system model. This preserves the differential algebraic structure of power networks. This also allows a seamless transition from reduced order to the full order dynamic and algebraic state variables.
  • Thorough time-domain simulations under various transient conditions have been conducted to evaluate the performance of the proposed techniques. The test system includes: (1) a modified IEEE 39-bus system that models detailed dynamics of conventional power plants, a solar plant operating in grid-forming mode, and algebraic constraint models, and (2) a 2000-bus Texas network incorporating comprehensive $11^{th}$-order models of conventional power plants. Additionally, to demonstrate the advantages of the proposed MOR techniques, a comparison with the commonly used linear ODE-based balanced truncation method is presented.
Paper Organization.   Sec. II presents the advanced power system model considered in this study. Sec. III delineates the problem formulation and scope. Sec. IV and V present the two MOR algorithms for NDAE power system models. Case studies are presented in Sec. VI while the paper is concluded in Sec. VII.
Notation.   Capital bold letters are used to represent matrices, as in $\bm{A}$ while small capital bold letters, such as $\bm{b}$, denote vectors. All the sets are represented in calligraphic fonts, such as $\mathcal{M}$ or $\mathcal{R}$. The symbol $\mathbb{R}^{u\times v}$ represents a real-valued matrix of size $u\times v$, similarly, $\mathbb{R}^{k}$ denotes a real-valued column vector with $k$ elements. The notations $\bm{O}$ and $\bm{I}$ denote zero and identity matrix of appropriate dimensions, respectively. The union of two sets is denoted by $\cup$. The symbol $\mathbb{S}^{n\times n}_{++}$ represents a square positive definite matrix of size $n\times n$ while the notation $\bm{e}_{j}$ represents a column vectors of zeros with $1^{\prime}s$ only at location $j$. Also, all quantities are given in per unit (p.u) unless otherwise specified.
II Nonlinear DAE Power System Model
We consider a grid model with $S$ solar power plants, $G$ conventional power plants, and $L_{k}$ loads. The overall power system is modeled as a graph with $\mathcal{N}=\left\{1,...,N\right\}$ as the set of nodes/buses and $\mathcal{E}$ as the set of edges or transmission lines. The set of buses are grouped into various types: =$\mathcal{R}=\left\{1,...,R\right\}$ represents buses with PV power plants, $\mathcal{G}=\left\{1,...,G\right\}$ denotes buses connected to the conventional power plants, $\mathcal{L}$ includes buses that contain loads, while $\mathcal{U}$ collects non-unit buses that are not connected to any elements.
@@ II Nonlinear DAE Power System Model
In model (1), the set of differential equations (1a) encompasses the dynamics of PV plants, conventional power plants, and composite loads dynamics, while the algebraic constraint (1b) models the current/power balance equations of the electrical network. The vector $\bm{x}_{d}(t)\in\mathbb{R}^{n_{d}}$ represents dynamic states and it lumps the dynamic variables of conventional power plants, PV plants, and composite load dynamics as $\bm{x}_{d}=\begin{bmatrix}\bm{x}_{G}^{\top}&\bm{x}_{R}^{\top}&\bm{x}_{L}^{\top}\end{bmatrix}^{\top}$ where $\bm{x}_{G}$ are the states of conventional power plants, $\bm{x}_{R}$ denotes the states of PV plants, and $\bm{x}_{L}$ represents the states of dynamic loads. The notation $\bm{x}_{a}(t)\in\mathbb{R}^{n_{a}}$ denotes algebraic states and it contains the states of the network (voltage and current phasors). The vector $\bm{w}(t)=\begin{bmatrix}\bm{P}_{d}^{\top}&\bm{I}_{s}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{w}}$ contains load demand $\bm{P}_{d}$ and sun irradiance $\bm{I}_{s}$ while $\bm{u}(t)=\begin{bmatrix}\bm{u}_{G}^{\top}&\bm{u}_{R}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{n_{u}}$ defines the system control inputs with $\bm{u}_{G}$ and $\bm{u}_{R}$ denoting the control inputs of generators and solar farms, respectively. Further detailed explanations of these vectors and complete dynamical equations (set of differential equations) describing the models of PV plants, conventional power plants, composite load dynamics, and system algebraic constraint model used in this study are given in Appendix A.
@@
where $\bm{E}\in\mathbb{R}^{n\times n}$ is a singular binary matrix encoding system algebraic model with rows of zeros, function ${\bm{f}}\left({\bm{x}},{\bm{u}},{\bm{w}}\right)$ represents the corresponding nonlinearity, $\bm{C}\in\mathbb{R}^{p\times n}$ is the output matrix, while the rest of the real-valued matrices ${\bm{B}}_{u}\in\mathbb{R}^{n\times n_{u}}$, ${\bm{A}}\in\mathbb{R}^{n\times n}$, ${\bm{B}}_{w}\in\mathbb{R}^{n\times n_{w}}$ maps the system control inputs $\bm{u}$, state vector $\bm{x}$, and the disturbance vector $\bm{w}$ in the power system dynamics. Furthermore, throughout the paper, we assume that in the NDAE model (2), the pair $(\bm{E},\bm{A})$ is regular, and the power system model is observable and controllable. These assumptions are common and power system NDAEs are known to be regular, controllable, and observable [2, 25].
@@ IV Structure-Preserving POD (SP-POD) MOR
3 Create snapshot matrices $\bm{X}_{d}$ (5a), $\bm{X}_{a}$ (5b), and $\bm{X}_{f}$ (10)
4 if $n<<t$ then
5      Perform SVD as:
7else
8      Perform eigenvalue decomposition:
10       Then design $\bm{W}_{d}$ and $\bm{W}_{a}$ as folows:
12Construct $\bm{W}=\mathrm{blkdiag}(\bm{W}_{d},\;\bm{W}_{a})$ as in (6)
14 Construct $\bm{T}=\mathrm{blkdiag}(\bm{T}_{d},\;\bm{T}_{a})$ as in (7)
15 Design $\bm{W}_{R}=\bm{T}\bm{W}$ and $\bm{W}_{L}=\bm{W}_{R}^{-1}$
16 SVD $\bm{X}_{f}$ (10) as: $\bm{X}_{f}=\bm{W}_{f}\bm{\Sigma}_{f}\bm{\Lambda}_{f}$
17 Start greedy algorithm to design $\bm{P}_{M}$
18 Construct rank-$p$ approximating basis
$\bm{W}_{fr}=\begin{bmatrix}\bm{w}_{f_{1}}\;,\bm{w}_{f_{2}}\;,\cdots,\bm{w}_{f_{p}}\end{bmatrix}$
19 Choose the first index: $[\rho$, $i_{1}]$ $=\mathrm{max}(\bm{w}_{f_{1}})$
20 Construct first measurement matrix
$\bm{P}_{M_{1}}=\bm{e}_{i_{1}},\,\bm{W}_{fr}=[\bm{w}_{f_{1}}]$
21 for $j=2:p$ do
22       calculate $\bm{c}$ using, $\bm{P}_{M}^{\top}\bm{W}_{fr}\bm{c}=\bm{P}_{M}\bm{w}_{f_{j}}$
23       compute residual, $\bm{d}=\bm{w}_{f_{j}}-\bm{W}_{fr}\bm{c}$
24       update $\bm{P}_{M}$ and $\bm{W}_{fr}$ as follows:
25       $[\rho,i_{j}]=\mathrm{max}(\bm{d})$
26       $\bm{W}_{fr}=[\bm{W}_{fr},\bm{w}_{f_{j}}],\,\bm{P}_{M}=[\bm{P}_{M},\bm{e}_{i_{j}}]$
Algorithm 1 SP-POD for power system NDAE models
@@ 26       $\bm{W}_{fr}=[\bm{W}_{fr},\bm{w}_{f_{j}}],\,\bm{P}_{M}=[\bm{P}_{M},\bm{e}_{i_{j}}]$
Therefore, we reduce the nonlinearity using the discrete empirical interpolation method (DEIM) [10, 26, 27]. The primary goal is to handle the nonlinear terms efficiently within a reduced-dimensional space instead of the full-dimensional space $\mathbb{R}^{n}$. DEIM achieves this by selecting a subset of spatial locations (interpolation points) where the nonlinear function is evaluated. By measuring specific points in the state space rather than the entire set of state variables, DEIM approximates the nonlinear term through interpolation around these selected points, this approximation can be expressed as [10]:
$\underbrace{\bm{W}_{L}\bm{W}_{fr}}_{r\times p}\underbrace{{\bm{f}}_{r}\left(\cdot\right)}_{p\times 1}.$ (9)
@@
To carry out such approximation for the nonlinear term we start by storing the snapshots of the ${\bm{f}}\left({{\bm{x}}},{\bm{u}},{\bm{w}}\right)$ in a matrix given as:
@@
Note that, to construct matrix $\bm{X}_{f}$, the function $\bm{f}(\bm{x},\bm{u},\bm{w})$ is evaluated for each time snapshot of $\bm{x}_{d}$, $\bm{x}_{a}$ (stored in data matrices $\bm{X}_{d}$ and $\bm{X}_{a}$), and the corresponding $\bm{u}$ and $\bm{w}$. For example $\bm{f}_{0}=\bm{f}(\bm{x}_{0},\bm{u}_{0},\bm{w}_{0}),\quad\bm{f}_{1}=\bm{f}(\bm{x}_{1},\bm{u}_{1},\bm{w}_{1})$, and so on. Then we take SVD of this snapshot matrix, $\bm{X}_{f}=\bm{W}_{f}\bm{\Sigma}_{f}\bm{\Lambda}_{f}$ with matrix $\bm{W}_{f}=\begin{bmatrix}\bm{w}_{f_{1}}\;,\bm{w}_{f_{2}}\;,\cdots,\bm{w}_{f_{n}}\end{bmatrix}\in\mathbb{R}^{n\times n}$ containing the left singular vectors, matrix $\bm{\Lambda}_{f}\in\mathbb{R}^{t\times n}$ lumping the right singular vectors, and $\bm{\Sigma}_{f}\in\mathbb{R}^{n\times n}$ being the diagonal matrix containing the HSVs. Next, we design $\bm{W}_{fr}=\begin{bmatrix}\bm{w}_{f_{1}}\;,\bm{w}_{f_{2}}\;,\cdots,\bm{w}_{f_{p}}\end{bmatrix}\in\mathbb{R}^{n\times p}$ as the first $p$ columns of $\bm{W}_{f}$. Finally, we design a binary measurement matrix $\bm{P}_{M}$ that selects optimal points in the reduced subspace $\bm{W}_{fr}$ so that nonlinearity can be reconstructed using the selected points efficiently. To construct such $\bm{P}_{M}$ we utilize residual-based greedy technique [10, 27], which essentially puts a measurement (or a $1$ in matrix $\bm{P}_{M}$) where residual/error is maximum (highlighting measurement point is required). These points are chosen to maximize the approximation accuracy. The greedy algorithm proceeds as follows:
@@
and the next measurement point is selected where the value of $\bm{d}$ is maximum. By selecting points with the maximum residual, the algorithm ensures that each new measurement location adds the most significant new information about the nonlinear term $\bm{f}(\cdot)$. After completion of the iterations, the approximation to the nonlinearity can be expressed as:
@@
We note that, if the approximation of the nonlinearity in the projected basis using the presented DIEM is insufficient, increasing the number of interpolation points can enhance accuracy. Also, if the computational cost of mapping the nonlinearity is low (such as form most of power system models), the DIEM step in the proposed algorithm may be omitted entirely. Additionally, other greedy algorithms such as Q-DIEM or EIM can also be utilized as effective substitutes for DIEM to efficiently reduce the computational complexity of handling the nonlinearity [10].
That said, the final ROM can be written as follows:
@@
The overall proposed SP-POD-based MOR algorithm is summarized in Algorithm 1. In the following section, we present the structure-preserving balanced POD-based (SP-BPOD) MOR technique for the complete NDAE representation of power systems.